OpenStudy (lgbasallote):

\[\LARGE \int \frac{(5-v)dv}{-2(v^2 - v - 2)}\]

OpenStudy (anonymous):

= -1/2 ∫ ( 5 - v) / ( v² -v -2) Adjust the numerator a little bit: = 1/4) ∫( 2v -1 ) / ( v² -v -2) dv - 9/4 ∫ dv/ ( v-2) ( v +1)

OpenStudy (lgbasallote):

would you happen to know how to do this via partial fractions?

OpenStudy (lgbasallote):

i think it's possible

OpenStudy (anonymous):

Split them out by numerator: = -5/2 ∫ dv/ ( v² -v -2) - ∫ v / ( v² -v -2)

OpenStudy (lgbasallote):

-5/2?

OpenStudy (anonymous):

Because when you split it the first part of numerator is 5

OpenStudy (lgbasallote):

oh lol yeah of course

OpenStudy (anonymous):

The second part: 1/2 ∫ v / ( v² -v -2)

OpenStudy (lgbasallote):

this is where i do the P.F.?

OpenStudy (anonymous):

Any method works fine :)

OpenStudy (lgbasallote):

v^2 - v - 2 = (v-2)(v+1) A(v+1) B(v-2) = v when v = 2 A(3) = 2 A = 2/3 when v = -1 B(-3) = -1 B = 1/3 \[\frac{v}{v^2 - v - 2} \implies \int \frac{2}{3(v-2)} + \int \frac{1}{3(v+1)}?\]

OpenStudy (anonymous):

Yep :)

OpenStudy (lgbasallote):

okay thanks :D i think i can go from here now

OpenStudy (anonymous):

I'm just back! Glad to help :)